OpenStudy (anonymous):

PLEASE HELP. WILL GIVE A MEDAL. Write a compound inequality to represent all of the numbers between -4 and 6. Write your answer using proper set notation.

jimthompson5910 (jim_thompson5910):

let's say x is your unknown number

jimthompson5910 (jim_thompson5910):

x is between -4 and 6 if we include both -4 and 6, this means x is either -4 or greater than -4, so, \(\Large x \ge -4\) AND x is either 6 or less than 6, so, \(\Large x \le 6\)

jimthompson5910 (jim_thompson5910):

if you combine \(\Large x \ge -4\) and \(\Large x \le 6\), you get \(\Large -4 \le x \le 6\)

jimthompson5910 (jim_thompson5910):

writing that in set notation will look like this \[\Large \left\{x|x\in\mathbb{R}, -4 \le x \le 6\right\}\] that basically says "x is a real number and it's between -4 and 6"

OpenStudy (anonymous):

Thank you so much. c:

jimthompson5910 (jim_thompson5910):

np